# A new Linear Time Bi-level $\ell_{1,\infty}$ projection ; Application to the sparsification of auto-encoders neural networks

### v1

@@ Abstract
In this paper, we propose a new bi-level projection method, for which we show that the time complexity for the $\ell_{1,\infty}$ norm is only $\mathcal{O}\big{(}nm\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$. Moreover, we provide a new $\ell_{1,\infty}$ identity with mathematical proof and experimental validation. Experiments show that our bi-level $\ell_{1,\infty}$ projection is $2.5$ times faster than the actual fastest algorithm and provides the best sparsity while keeping the same accuracy in classification applications.
Keywords— Bi-level Projection, Structured sparsity, low computational linear complexity
1 Introduction
@@
The paper is organized as follows. First, we provide the current state of the art of the $\ell_{1,\infty}$ ball projection. Then, we provide in section 3 the new bi-level $\ell_{1,\infty}$ projection. In section 4, we apply our bi-level framework to other constraints providing sparsity, such as $\ell_{1,1}$ and $\ell_{1,2}$ constraints. In Section 5, we finally compare different projection methods experimentally. First, we provide an experimental analysis of the projection algorithms onto the bi-level projection $\ell_{1,\infty}$ ball. This section shows the benefit of the proposed method, especially for time processing and sparsity. Second, we apply our framework to the classification using a supervised autoencoder on two synthetic datasets and a biological dataset.
2 State of the art of the $\ell_{1,\infty}$ ball projection
@@
$\mathcal{B}^{\eta}_{1,\infty}:=\left\{X\in\mathbb{R}^{n\times m}\ :\ {\|{X}\|}_{1\infty}\leq\eta\right\}.$ (2)
The projection $P_{\mathcal{B}^{\eta}_{1,\infty}}$ onto $\mathcal{B}^{\eta}_{1,\infty}$ , also noted $P^{1,\infty}_{\eta}$ in the sequel, is given by:
$P^{1,\infty}_{\eta}(Y)~{}=~{}P_{\mathcal{B}^{\eta}_{1,\infty}}(Y)~{}=~{}{arg\min\limits_{X\in\mathcal{B}^{\eta}_{1,\infty}}\frac{1}{2}\|X-Y\|_{\mathrm{F}}^{2}}$ (3)
@@
$P_{\mathcal{B}^{\alpha}_{1,\infty}}(Y)=Y-\mathrm{prox}_{\alpha\|\cdot\|_{\infty,1}}(Y)$ (6)
@@
3 A new Bi-level $\ell_{1,\infty}$ structured projection
3.1 A new bi-level projection
@@
Remark 3.2.
We say that $BP^{1,\infty}_{\eta}$ is a clipping operator, and $\hat{u}$ is its clipping threshold.
@@
and then with remark 3.1
@@
3.2 The $\ell_{1,\infty}$ identity
In the case of $\ell_{1,\infty}$ projection, we needed Moreau’s identity to develop the projection algorithm from the ”Prox”. In the case of our new bilevel $\ell_{1,\infty}$ projection, we have a direct linear-complexity algorithm that does not require Moreau’s identity. The aim of this section is to show the respective properties of these two projections. We study a norm of the projected regularized solution versus a norm of the corresponding residual [33]. Recall the classical triangle inequality, which is a consequence of the Cauchy–Schwartz inequality,
@@ 3.2 The $\ell_{1,\infty}$ identity
${\|{Y-BP^{1,\infty}_{\eta}(Y)}\|}_{1\infty}+{\|{BP^{1,\infty}_{\eta}(Y)}\|}_{1\infty}={\|{Y}\|}_{1\infty}\$ (18)
@@
Remark 3.4.
Careful examination of the $P^{1,\infty}_{\eta}$ projector algorithm [30, 23, 22] shows that $P^{1,\infty}_{\eta}$ is also a clipping operator, for a different threshold $u$ (See Line 15 of algorithm 1 in [30]) and thus projection verifies 16.
Proposition 3.5.
@@
${\|{Y-P^{1,\infty}_{\eta}(Y)}\|}_{1\infty}+{\|{P^{1,\infty}_{\eta}(Y)}\|}_{1\infty}={\|{Y}\|}_{1\infty}\$ (19)
The proof of Eq 19 follows the same way as for Eq 18. In fact, identities such as 18 and 19 hold for infinitely many clipping operators. A vector $u$ is a feasible clipping threshold if it satisfies bounds of remark 3.1 and sum to $\eta$.
Remark 3.6.
@@
3.3 Convergence and Computational complexity
@@
4 Extension to other sparse structured projections
@@
In this section we extend our bilevel method to the $\ell_{1,1}$ and $\ell_{1,2}$ balls, yielding structured sparsity.
4.1 Bilevel $\ell_{1,1}$ projection
@@
Let $x_{j}~{}=~{}P_{u_{j}}^{1}(y_{j})$, there exists some positive $\lambda_{j}$ verifying a critical equation, such that (section 6.5.2 of [35]):
$x_{j}=(y_{j}-\lambda_{j})_{+}-(-y_{j}-\lambda_{j})_{+}$ and $|x_{j}|_{1}=u_{j}$
@@
$|Y_{i,j}~{}-~{}X_{i,j}|_{1}$ $=\sum_{i}|~{}|Y_{i,j}|-\max(|Y_{i,j}|-\lambda_{j},0)|$ (22)
$=\sum_{i}|Y_{i,j}|-\sum_{i}\max(|Y_{i,j}|-\lambda_{j},0)$
which can be written as
@@
5 Experimental results
5.1 Benchmark times using PyTorch C++ extension using a MacBook Laptop with an i9 processor; Comparison with the best actual projection method
The experiments were run on a laptop with a I9 processor having 32 GB of memory. The state of the art on such is pretty large, starting with [22] who proposed the first algorithm, the Newton-based root-finding method and column elimination method [24, 23], and the recent paper of Chu et al. [25] which outperforms all the other state-of-the-art methods. We compare our bi-level method against the best actual algorithm proposed by Chu et al. which uses a semi-smooth Newton algorithm for the projection. We use C++ implementation provided by the authors and the PyTorch C++ implementation of our bi-level $\ell_{1,\infty}$ method is based on fast $\ell_{1}$ projection algorithms of [20, 21] which are of linear complexity. The code is available online [^1] https://github.com/memo-p/projection.
Figure 1: Comparison of our bi-level method with Chu et al. Processing time as a function of the radius (matrices fixed, 1000x10000)
The size of the matrices is, 1000x10000, values between 0 and 1 uniformly sampled. Figure 1 shows that the running time as a function of the radius is at least 2.5 times faster than the actual fastest method, Chu et al.. Note that the running time of the bi-level one is almost not impacted by the radius.
Figure 2: Processing time using C++: our bi-level projection method versus Chu et. al. method. m=1000 and $\eta=1$ fixed .
Figure 2 shows the running time as a function of the matrix size. Here the radius has been fixed to $\eta=1$. The running time of our bilevel method is at least 2.5 faster than that of the actual fastest method when increasing the matrix size both in number of columns and rows. Note that PyTorch c++ extension is 20 times faster than the standard PyTorch implementation.
5.2 Benchmark of Identity Proposition
We generate two artificial biological datasets to benchmark our bi level projection using the $make\_classification$ utility from scikit-learn. We generate $n=1,000$ samples with a number of $m=1000$ features. The first one with 64 (data-64) informative features and the second (data-16) with 16 informative features
@@ 5.2 Benchmark of Identity Proposition
Figure 3: Identity norm comparison of our bi-level $\ell_{1,\infty}$ projection method versus classical.
Fig 3 shows that the two curves ( Bilevel and usual $\ell_{1,\infty}$ projections) and the $\eta$ are parameter are perfectly coincident, and perfectly linear as expected by the identity equations [18].
Figure 4: Identity norm $L_{2}$ error comparison of our bi-level projection method versus classical.
Fig 4 shows that $P\ell_{1,\infty}$ has the best $L_{2}$ error, However $L_{2}$ error is not more relevant for our purpose than any other norm.
@@
Fig 5 shows that the curve Bilevel $\ell_{1,1}$ is perfectly linear as expected by the identity equation 24.
Figure 6: Identity norm, our bi-level $\ell_{1,2}$ projection method.
Fig 6 shows that the curve Bilevel $\ell_{1,2}$ projection satisfies the identity matrix relation.
Figure 7: 64 informative features Sparsity Top the Bilevel $\ell_{1,\infty}$, Middle: Bilevel $\ell_{1,1}$, bottom: Bilevel $\ell_{1,2}$ projection
Figure 8: 16 informative features. Sparsity Top the Bilevel $\ell_{1,\infty}$, Middle: Bilevel $\ell_{1,1}$, bottom: Bilevel $\ell_{1,2}$ projection
Cumul Sparsity bi-level $\ell_{1,\infty}$ bi-level $\ell_{1,1}$ bi-level $\ell_{1,\infty}$ $\ell_{1,\infty}$
@@
Table 1: Comparison of Sparsity for two datasets with different informative features
Table 1 shows that our bilevel $\ell_{1,\infty}$ projection outperforms sparsity of bilevel $\ell_{1,1}$ projection and bilevel $\ell_{1,2}$. Although the sparsity curves look very close; Table 1 shows that bilevel $\ell_{1,1}$ projection is slightly more sparse than bilevel $\ell_{1,2}$.
@@
5.3 Experimental results on classification and feature selection using a supervised autoencoder neural network
5.3.1 Supervised Autoencoder (SAE) framework
Autoencoders were introduced within the field of neural networks decades ago, their most efficient application being dimensionality reduction [36, 37]. Autoencoders were used in application ranging from unsupervised deep-clustering [38, 39] to supervised learning adding a classification loss in order to improve classification performance [40, 41]. In this paper, we use the cross entropy as the added classification loss.
Let $X$ be the dataset in $\mathbb{R}^{d}$, and $Y$ the labels in $\{0,\dots,k\}$, with $k$ the number of classes. Let $Z\in\mathbb{R}^{k}$ be the encoded latent vectors, $\widehat{X}$ $\in\mathbb{R}^{d}$ the reconstructed data and $W$ the weights of the neural network. Note that the dimension of the latent space $k$ corresponds to the number of classes.
@@ 5.3.1 Supervised Autoencoder (SAE) framework
where $\phi(X,Y)=$ $\alpha\psi(X,\widehat{X}))+$ $\mathcal{H}(Y,Z)$. We use the Cross Entropy $\mathcal{H}$ as the added classification loss and the robust Smooth $\ell_{1}$ (Huber) Loss [42] as the reconstruction loss $\psi$. Parameter $\alpha$ is a linear combination factor used to define the final loss. We compute the mask by using the various bilevel projection methods, and we use the double descent algorithm [43, 44] for minimizing the criterion 28. We implemented our SAE method using the PyTorch framework for the model, optimizer, schedulers and loss functions. We use a fully connected neural network with only one hidden layer (dimension 100) and a latent layer of dimension $k=2$ We chose the ADAM optimizer [45], as the standard optimizer in PyTorch. We used a symmetric linear fully connected network with the encoder comprised of an input layer of $d$ neurons, one hidden layer followed by a ReLU or SiLU activation function.
5.3.2 Experimental accuracy results on autoencoder neural neworks
Figure 9: Accuracy as a function of the radius parameter $\eta$; Left, 64 informative features :16 informative features
Figure 9 shows the impact of the radius ($\eta$) on a synthetic dataset using by the bilevel projection. We consider now one of the real dataset from the study of in single-cell CRISPRi screening [46]. This dataset is composed of 779 cells and 10,000 features (genes).

### v2

@@ Abstract
In this paper, we propose a new bi-level projection method, for which we show that the time complexity for the $\ell_{1,\infty}$ norm is only $\mathcal{O}\big{(}nm\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$.
Moreover, we provide a new $\ell_{1,\infty}$ identity with mathematical proof and experimental validation.
Experiments show that our bi-level $\ell_{1,\infty}$ projection is $\mathcal{O}\big{(}\log(nm)\big{)}$ times faster than the actual fastest algorithm. Our bi-level $\ell_{1,\infty}$ projection outperforms the sparsity of the usual $\ell_{1,\infty}$ projection while keeping the same or slightly better accuracy in classification applications.
Index Terms:
Bilevel Projection, Structured sparsity, low computational linear complexity
I Introduction
@@
The paper is organized as follows. First, we provide the current state of the art of the $\ell_{1,\infty}$ ball projection. Then, we provide in section III the new bi-level $\ell_{1,\infty}$ projection. In section IV, we apply our bi-level framework to other constraints providing sparsity, such as $\ell_{1,1}$ and $\ell_{1,2}$ constraints. In Section V, we finally compare different projection methods experimentally. First, we provide an experimental analysis of the projection algorithms onto the bi-level projection $\ell_{1,\infty}$ ball. This section shows the benefit of the proposed method, especially for time processing and sparsity. Second, we apply our framework to the classification using a supervised autoencoder on two synthetic datasets and a biological dataset.
II State of the art of the $\ell_{1,\infty}$ ball projection
@@
$\mathcal{B}^{1,\infty}_{\eta}:=\left\{X\in\mathbb{R}^{n\times m}\ :\ {\|{X}\|}_{1,\infty}\leq\eta\right\}.$ (2)
The projection $P_{\mathcal{B}^{1,\infty}_{\eta}}$ onto $\mathcal{B}^{1,\infty}_{\eta}$ , also noted $P^{1,\infty}_{\eta}$ in the sequel, is given by:
$P^{1,\infty}_{\eta}(Y)~{}=~{}P_{\mathcal{B}^{1,\infty}_{\eta}}(Y)~{}=~{}{arg\min\limits_{X\in\mathcal{B}^{1,\infty}_{\eta}}\frac{1}{2}\|X-Y\|_{\mathrm{F}}^{2}}$ (3)
@@
$P_{\mathcal{B}^{{1,\infty}}_{\alpha}}(Y)=Y-\mathrm{prox}_{\alpha\|\cdot\|_{\infty,1}}(Y)$ (6)
@@
III A new Bi-level $\ell_{1,\infty}$ structured projection
III-A A new bi-level projection
@@
Remark III.2.
We say that $Y\rightarrow BP^{1,\infty}_{\eta}$ is a clipping operator, and $\hat{u}$ is its clipping threshold.
@@
and then with remark III.1
@@
III-B The $\ell_{1,\infty}$ identity
In the case of $\ell_{1,\infty}$ projection, we needed Moreau’s identity to develop the projection algorithm from the "Prox". In the case of our new bilevel $\ell_{1,\infty}$ projection, we have a direct linear-complexity algorithm that does not require Moreau’s identity. The aim of this section is to show the respective properties of these two projections. We study a norm of the projected regularized solution versus a norm of the corresponding residual [33]. Recall the classical triangle inequality, which is a consequence of the Cauchy–Schwartz inequality,
@@ III-B The $\ell_{1,\infty}$ identity
${\|{Y-BP^{1,\infty}_{\eta}(Y)}\|}_{1,\infty}+{\|{BP^{1,\infty}_{\eta}(Y)}\|}_{1,\infty}={\|{Y}\|}_{1,\infty}\$ (18)
@@
Remark III.4.
Careful examination of the $P^{1,\infty}_{\eta}$ projector algorithm [30, 23, 22] shows that $P^{1,\infty}_{\eta}$ is also obtained by a clipping operator, for a different threshold $u$ (See Line 15 of algorithm 1 in [30]) and thus projection verifies 16.
Proposition III.5.
@@
${\|{Y-P^{1,\infty}_{\eta}(Y)}\|}_{1,\infty}+{\|{P^{1,\infty}_{\eta}(Y)}\|}_{1,\infty}={\|{Y}\|}_{1,\infty}\$ (19)
The proof of Eq 19 follows the same way as for Eq 18. In fact, identities such as 18 and 19 hold for infinitely many clipping operators. A vector $u$ is a feasible clipping threshold if it satisfies bounds of remark III.1 and sum to $\eta$.
Remark III.6.
@@
III-C Convergence and Computational complexity
@@
IV Extension to other sparse structured projections
@@
In this section, we extend our bilevel method to the $\ell_{1,1}$ and $\ell_{1,2}$ balls, yielding structured sparsity.
IV-A Bilevel $\ell_{1,1}$ projection
@@
The projection $x_{j}~{}=~{}P_{u_{j}}^{1}(y_{j})$ is obtained by elementwise soft thresholding (proposition 2.2 in [20] or section 6.5.2 in [35]), that is, there exists some positive unique $\lambda_{j}$ verifying a critical equation such that:
$x_{j}=\max(y_{j}-\lambda_{j},0)-\max(-y_{j}-\lambda_{j},0)$
and
$|x_{j}|_{1}=u_{j}$
so
@@
$|Y_{i,j}~{}-~{}X_{i,j}|$ $=|~{}|Y_{i,j}|-\max(|Y_{i,j}|-\lambda_{j},0)~{}|$ (22)
$=|Y_{i,j}|-\max(|Y_{i,j}|-\lambda_{j},0)$
$=|Y_{i,j}|-|X_{i,j}|$
which can be written, by summing on $i$:
@@
V Experimental results
V-A Benchmark times using PyTorch C++ extension using a MacBook Laptop with an i9 processor; Comparison with the best actual projection method
The experiments were run on a laptop with a I9 processor having 32 GB of memory. The state of the art on such is pretty large, starting with [22] who proposed the first algorithm, the Newton-based root-finding method and column elimination method [24, 23], and the recent paper of Chu et al. [25] which outperforms all the other state-of-the-art methods. We compared our bi-level method against all the existing projection algorithms, and our algorithm is faster in all the scenarios. The algorithm from [22] is on average 30 times slower than our algorithm, note that this factor growth logarithmically with data size. As shown in [25], the best actual algorithm is proposed by Chu et al. which uses a semi-smooth Newton algorithm for the projection. That is why we focus our presentation on comparing against this particular method. We use the C++ implementation provided by the authors and the PyTorch C++ implementation of our bi-level $\ell_{1,\infty}$ method is based on fast $\ell_{1}$ projection algorithms of [20, 21] which are of linear complexity. The code of all the compared algorithms is available online [^1] https://github.com/memo-p/projection.
! Refer to caption Refer to caption
Figure 1: Processing time using C++ as a function of the number of features $n=1000$ samples (top) and Samples $m=1000$ features (bottom): bi-level projection method versus Chu et al. method.
Figure 1 shows the running time as a function of the matrix size. Here the radius has been fixed to $\eta=1$. The fitting of a linear curve (red) on the data (blue) shows that the running time of our bilevel $\ell_{1,\infty}$ projection is linear with the number of features and the number of samples. The fitting of a $nlogn$ curve (green) on the data (orange) shows that the running time of the usual $\ell_{1,\infty}$ projection grows as $nlogn$ with the number of features and the number of samples. Moreover, the slope of the usual projection algorithm is greater by a factor 2.5 on both graphs than the slope of the bilevel algorithm.
Figure 2: Processing time using C++ as a function of the number of features (Top), and samples (bottom)
Figure 2 shows that all bilevel algorithms have the same slopes for time versus feature or sample number.
Note that PyTorch c++ extension is 20 times faster than the standard PyTorch implementation.
V-B Benchmark of Identity Proposition
We generate two artificial biological datasets to benchmark our bi level projection using the $make\_classification$ utility from scikit-learn. We generate $n=1,000$ samples with a number of $m=1,000$ features. The first one with 64 (data-64) informative features and the second (data-16) with 16 informative features
@@ V-B Benchmark of Identity Proposition
Figure 3: Identity norm comparison Top: the Bilevel $\ell_{1,\infty}$ versus classical, Middle: Bilevel $\ell_{1,1}$, bottom: Bilevel $\ell_{1,2}$ projection.
Fig 3 shows that the two curves (Bilevel and usual $\ell_{1,\infty}$ projections) and the $\eta$ parameter are perfectly coincident, and perfectly linear as expected by the identity equation (18).
Figure 4: Bilevel $\ell_{1,\infty}$ projection and usual $\ell_{1,\infty}$ projection with $\ell_{2,2}$ norm.
Remark V.1.
The identity equations hold only if the norm is similar to the projection. Figure 4 shows that the identity equation is not true when using Bilevel and usual $\ell_{1,\infty}$ projections and the classical $\ell_{2,2}$ norm. Usual projection $\ell_{1,\infty}$ has the lower $\ell_{2,2}$ error. However, $\ell_{2,2}$ error is not more relevant for our purpose than any other norm.
Figure 5: 64 informative features Sparsity Top: the Bilevel $\ell_{1,\infty}$, Middle: Bilevel $\ell_{1,1}$, bottom: Bilevel $\ell_{1,2}$ projection
Figure 6: 16 informative features. Sparsity: Top the Bilevel $\ell_{1,\infty}$, Middle: Bilevel $\ell_{1,1}$, bottom: Bilevel $\ell_{1,2}$ projection
@@
[…]
Table I shows that our bilevel $\ell_{1,\infty}$ projection outperforms sparsity of the usual $\ell_{1,\infty}$ projection. Although the sparsity curves look very close. Table I shows that bilevel $\ell_{1,1}$ projection is slightly more sparse than bilevel $\ell_{1,2}$.
@@
V-C Experimental results on classification and feature selection using a supervised autoencoder neural network
V-C1 Supervised Autoencoder (SAE) framework
Autoencoders were introduced within the field of neural networks decades ago, their most efficient application being dimensionality reduction [36]. Autoencoders were used in application ranging from unsupervised deep-clustering [37, 38] to supervised learning, adding a classification loss in order to improve classification performance [39, 40]. In this paper, we use a supervised autoencoder with the cross entropy as the added classification loss.
Let $X$ be the concatenated raw data matrix ($n\times m$) (n is the number of samples (cells) and m the number of genes). Let $\widehat{X}$ the reconstructed data and $W$ the weights of the neural network. Let $Z$ the encoded latent space. Note that the dimension of the latent space $k$ corresponds to the number of classes.
@@ V-C1 Supervised Autoencoder (SAE) framework
where $\phi(X,Y)=$ $\alpha\cdot\psi(X,\widehat{X})+$ $\mathcal{H}(Y,Z)$. We use the robust Smooth $\ell_{1}$ (Huber) Loss [41] as the reconstruction loss $\psi$. Parameter $\alpha$ is a linear combination factor used to define the final loss. We compute the mask by using the various bilevel projection methods, and we use the double descent algorithm [42, 43] for minimizing the criterion 28. We implemented our SAE method using the PyTorch framework for the model, optimizer, schedulers and loss functions. We use a fully connected neural network with only one hidden layer (dimension 100) and a latent layer of dimension $k=2$. We chose the ADAM optimizer [44], as the standard optimizer in PyTorch. We use the smooth SiLU activation function.
V-C2 Experimental accuracy results on autoencoder neural networks
Figure 7: Accuracy as a function of the radius parameter $\eta$; Top, 64 informative features :Bottom 16 informative features
Figure 7 shows the impact of the radius ($\eta$) on a synthetic dataset using the bilevel projection versus the usual projection. It can be seen that the best accuracy is obtained for $\eta=0.5$ for $\ell_{1,\infty}$ and for $\eta=1$ for the bilevel $\ell_{1,\infty}$ projection. Moreover, the accuracy is more robust to parameter $\eta$ for the bilevel projection. Table II shows accuracy classification for 64 informative features. The baseline is an implementation that does not process any projection. Compared to the baseline, the SAE using the $\ell_{1,\infty}$ projection improves the accuracy by $10.3\%$.
Table III shows that accuracy results for 16 informative features of the bilevel $\ell_{1,\infty}$ and classical $\ell_{1,\infty}$ is slightly better.
From Tables II on synthetic dataset it can be seen that the best accuracy is obtained for $\eta=0.5$ for $\ell_{1,\infty}$ and for $\eta=1.$ for the bilevel $\ell_{1,\infty}$ projection. Maximum accuracy of both method are similar. However, sparsity and computation time are better for bilevel $\ell_{1,\infty}$ than for regular $\ell_{1,\infty}$.
Synthetic 64 Baseline $\ell_{1,\infty}$ bilevel $\ell_{1,\infty}$
Best Radius - 1 2.0
[…]
TABLE II: Synthetic dataset 64 features. SiLU activation, Accuracy : comparison of $\ell_{1,\infty}$ and bi-level $\ell_{1,\infty}$.
Synthetic 16 Baseline $\ell_{1,\infty}$ bilevel $\ell_{1,\infty}$
Best Radius - 0.5 1.0
[…]
TABLE III: Synthetic dataset 16 features. SiLU activation, Accuracy : comparison of $\ell_{1,\infty}$ and bi-level $\ell_{1,\infty}$.
We consider the HIF2 dataset, now one of the real dataset from the study of in single-cell CRISPRi screening [45]. This HIF2 dataset is composed of 779 cells and 10,000 features (genes) [45].